OpenStudy (anonymous):

Please Help!!! I will give medals and I will fan!!!!!!!!!!!!!! Simplify: (4a + 2b)(a – b) A. 4a2– 6ab – 2b2 B. 4a2 – 2ab – 2b2 C. 4a2 – 8ab – 2b2 D. 4a2 + 2ab – 2b2

OpenStudy (anonymous):

is itA.

OpenStudy (anonymous):

Yes it is A....

OpenStudy (anonymous):

can u help me with more

OpenStudy (anonymous):

As (4a+2b)a-(4a+2b)b so =4a^2+2ab-4ab-2b^2 then =4a^2-2b^2-2ab

OpenStudy (anonymous):

B is right.

OpenStudy (anonymous):

Sorry, my keyboard was broken.

OpenStudy (anonymous):

I can help you with more.

OpenStudy (anonymous):

Simplify. (4a – 3)^2 A. 162a – 12 + 9 B. 8a2 – 24a + 6 C. 16a2 – 9 D. 16a2 – 24a + 9

OpenStudy (anonymous):

(4a – 3)^2 means (4a – 3)(4a – 3).

OpenStudy (anonymous):

16a^2-9 is that right

OpenStudy (anonymous):

Let me see. (4a – 3)(4a – 3) =(4a-3)4a-(4a – 3)3 =16a^2-12a-12a+9 So the answer is 16a^2-24a+9

OpenStudy (anonymous):

Simplify. (x – 4)(x2 – 3x – 9) A. x3 – 4x2 – 12x + 36 B. x3 + x2 – 3x – 36 C. x3 – 7x2 – 21x + 36 D. x3 – 7x2 + 3x + 36

OpenStudy (anonymous):

You should try to do it yourself, and then I will check your work.

OpenStudy (anonymous):

I got D.

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

Great, you are right!

OpenStudy (anonymous):

Simplify (x2 + 5x + 12)(x2 + 6x + 7). A. x4 + 37x2 + 84 B. x4 + 30x2 + 84 C. x4+ 11x3 + 49x2 + 107x + 84 D. x4 + 6x3 + 7x2 + 11x + 19 i tried but i cant get it

OpenStudy (anonymous):

i got D.

OpenStudy (anonymous):

All right let me help you.

OpenStudy (anonymous):

I think the right answer is C.

OpenStudy (anonymous):

i have one more quetion

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Find the volume of the box. Use the formula V = lwh.  A. 6x3 + 13x2 + x – 2 B. 6x3+ 6x – 2 C. 6x3 – 2 D. 6x2 – 17x2 + x – 2